OpenStudy (anonymous):

Match the equation with its graph. -6/7x-1/2y=3/7 Just give me a rough draft of the graph

OpenStudy (anonymous):

*Please :)

hero (hero):

I'm pretty sure you mean this: \[-\frac{6}{7}x -\frac{1}{2}y = \frac{3}{7}\] But if you are going to post your fractions linearly like that at least use parentheses: -(6/7)x-(1/2)y=3/7 This lets us know that x and y are being multiplied by the fraction. If you leave it posted the way you have, it is ambiguous and someone could interpret x and y as being in the denominator of the fraction.

OpenStudy (anonymous):

Yes I did thanks uhm what would the graph kinda look like?

hero (hero):

First you have to put the equation in slope-intercept form. After that, then you can plug it in to your graphing calculator to the graph.

OpenStudy (anonymous):

slope is y=mx+b right?

OpenStudy (anonymous):

and isn't my slope 3/7?

hero (hero):

slope- `intercept form` is y = mx + b and 3/7 is `not` the slope

OpenStudy (anonymous):

what is y? in that equation is it 7 or 3? I was never much good at slope

hero (hero):

You will have to do some algebraic manipulation to convert the equation to the slope-intercept form.

OpenStudy (anonymous):

nope it would be -1 -1=m(12)+b

OpenStudy (anonymous):

so would slope be -1/12? I know i have to find m and b so no that wouldn't be slope